Two boats, A and B, move away from a buoy anchored at the middle of a river along the mutually perpendicular straight lines: the boat A along the river, and the boat B across the river. Having moved off an equal distance from the buoy the boats returned. Find the ratio of times of motion of boats τA/τB if the velocity of each boat with respect to water is η = 1.4 times greater than the stream velocity.
1
Expert's answer
2020-10-15T10:42:00-0400

Solution

Let boat A covered "l" distance, "v" is stream velocity, "v '" is each boat velocity with respect to water.

So time is taken by Boat A is given by

"t_A=\frac{l}{v'+v}+\frac{l}{v'-v}"

Similarly for Boat B is

"t_B=\frac{l}{\sqrt{v'^2-v^2}}+\frac{l}{\sqrt{v'^2-v^2}}=\frac{2l}{\sqrt{v'^2-v^2}}"

Therefore

Ratio of time taken by Boats is given by

"\frac{t_A}{t_B}=\frac{v'}{\sqrt{v'^2-v^2}}=\frac{\eta}{\sqrt{\eta^2-1}}"

Where "\eta=\frac{v'}{v}=1.4"

Therefore ratio becomes

"\frac{t_A}{t_B}=\frac{1.4}{\sqrt{(1.4)^2-1}}=1.42"

Therefore ratio of time taken by both boats is 1.42